# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v5

@@ IV-D Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[tim1]. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 8. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
@@
Figure 9: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance. Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
Figure 10: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ V-C Implementation Details
To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
@@ V-D2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
@@ V-D3 Results on FoggyTusimple
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83

### v6

@@ IV-D Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 8. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane[32] IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN[21] IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN[21] IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
@@
Figure 9: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane[32] IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN[21] IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN[21] IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
Figure 10: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ V-C Implementation Details
To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39, 32, 21], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
@@ V-D2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
@@ V-D3 Results on FoggyTusimple
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane[32] IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN[21] IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN[21] IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83